OpenStudy (anonymous):

verify that f(x) and G(x) are inverse functions. f(x) = -9x + 10, g(x) = -1/9x + 10/9

OpenStudy (anonymous):

@PrinceT94 could you help again?

OpenStudy (anonymous):

Solve for f(g(x)) and g(f(x)) and if they equal to x, they are inverse functions.

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

f(g(x)) = -9(-1/9x + 10/9) + 10 = x -10 +10 = x g(f(x) = -1/9(-9x + 10) + 10/9 = x -10/9 + 10/9 = x They both result in x, therefore they are inverse functions. Hope that helped!